### v3

# On The Collatz Conjecture: Topological and Ergodic Approach

Abstract.
We study the Collatz function famously related to the Collatz Conjecture under the topological and ergodic perspectives, including an approach with thermodynamic formalism. By introducing a key topology and its Borel $\sigma$-algebra we show that recurrence implies periodicity. Moreover, we establish that the set of periodic orbits is finite if, and only if, every continuous potential possesses some equilibrium state. The uniqueness of periodic orbits is equivalent to the uniqueness of equilibrium state for every bounded and continuous potential. Additionally, by using the Alexandroff compactification of the established topology, we prove finiteness of cycles, which is a significant advance to the conjecture itself. We also present a proof of uniqueness of cycle and show that there is no divergent orbits. Finally, we apply our technique to the Baker and Syracuse maps, obtaining a similar result on the finiteness of orbits and no divergence for a general class of important maps.
1. Introduction
The famous Collatz function $f:\mathbb{N}\to\mathbb{N}$ is defined as follows:
[…]
It has been extensively studied because of the famous Collatz Conjecture: For each $n\in\mathbb{N}$ there exists $k\in\mathbb{N}$ such that $f^{k}(n)=1$. In other words, every orbit enters the cycle $\{1,2,4\}$. This is the unique cycle known.
We intend to obtain some results on the topological and ergodic aspects of the Collatz function. It will be useful in the study of the Collatz Conjecture. We highlight its difficult by building a bridge between it and the existence and uniqueness of equilibrium states for every continuous potentials with respect to a key topology. It provides a ”dictionary” which perhaps turns possible the complete proof in another realm. We provide a significant advance by proving finiteness of cycles and later on, a proof of uniqueness of cycle. To prove the conjecture completely, we also prove that we cannot have divergent orbits. We also apply our technique for the Baker and Syracuse maps.
2. Main result
The following result is an attempt to link a very hard and famous problem on Number Theory to Ergodic Theory in the realm of Themodynamic Formalism. We list some topological and ergodic properties of the Collatz function which highlights both the beauty and the difficulty of the Collatz Conjecture using the language of equilibrium states.
Theorem A.
We have the following facts with respect to $f$
  • There exists a topology coarser than the discrete one and a $\sigma$-algebra with respect to which every recurrent point is periodic and the ergodic probabilities are the ones supported on the periodic orbits.
  • $f$ is measurable with respect to this $\sigma$-algebra but not continuous, once $f$ is only continuous with respect to the discrete one.
  • Every periodic orbit is an open subset.
  • Every $f$-invariant probability is a convex linear combination of ergodic probabilities and has zero entropy.
  • Finiteness of periodic orbits is guaranteed by the existence of an equilibrium state for every continuous potential $\phi:\mathbb{N}\to\mathbb{N}$.
  • Uniqueness of periodic orbits is guaranteed by the uniqueness of equilibrium state for every bounded and continuous potential $\phi:\mathbb{N}\to\mathbb{N}$.
3. Proof of Theorem A
We divide the proof of our Main Theorem into some lemmas and remarks. We proceed with this now.
3.1. Invoking foundations
The first lemma invokes foundations of General Topology and Measure Theory to build both a key topology and a key $\sigma$-algebra.
Lemma 1.
Let $\mathcal{T}_{\lambda}$ be a family of topologies on $\mathbb{N}$ and $\sigma(\mathcal{T}_{\lambda})$ their Borel $\sigma$-algebras. We have that
$\mathcal{T}:=\bigcap_{\lambda}\mathcal{T}_{\lambda}\quad\text{and}\quad\Sigma:=\bigcap_{\lambda}\sigma(\mathcal{T}_{\lambda})$
are respectively a topology and a $\sigma$-algebra on $\mathbb{N}$ if a map $f:\mathbb{N}\to\mathbb{N}$ is measurable with respect to every $\sigma(\mathcal{T}_{\lambda})$ it does with respect to $\Sigma$.
Proof.
It is well known from the theory of General Topology and Measure Theory that $\mathcal{T}$ and $\Sigma$ are respectively a topology and a $\sigma$-algebra. Moreover, given any $A\in\Sigma$ we have $A\in\sigma(T_{\lambda})$ for every $\lambda$, which implies $f^{-1}(A)\in\sigma(T_{\lambda})$ for every $\lambda$ because $f$ is measurable with respect to $\sigma(\mathcal{T}_{\lambda})$ and we conclude that $f^{-1}(A)\in\Sigma$. Since $A$ is arbitrary, we obtain $f$ measurable with respect to $\Sigma$. ∎
3.2. Trapping the orbits towards periodicity
The following lemma builds both a topology and its Borel $\sigma$-algebra making the Collatz map both measurable and predictable in the language of Ergodic Theory.
Lemma 2.
There exist a topology $\mathcal{T}$ and a Borel $\sigma$-algebra $\Sigma$ with respect to which the Collatz map $f:\mathbb{N}\to\mathbb{N}$ possesses $f$-invariant Borel probabilities. Moreover, every recurrent point is periodic.
Proof.
Endow $\mathbb{N}$ with $\mathcal{T}$ as the intersection of all topologies (the coarsest) containing the following collection of subsets
$\{\{n,2n\}\mid n\in\mathbb{N}\}.$
By Lemma 1 $\mathcal{T}$ is well defined and we consider the $\sigma$-algebra $\Sigma$ constructed also as in Lemma 1. We obtain $f$ measurable with respect to $\Sigma$.
Let $\mathcal{M}_{f}(\mathbb{N})$ be the set of all Borelian $f$-invariant probabilities. Once there exists at least one periodic orbit, that is, $1\to 4\to 2\to 1$, we have $\mathcal{M}_{f}(\mathbb{N})\neq\emptyset$ because we can take the ergodic probability $\delta_{0}$ supported at this orbit.
Take any $\mu\in\mathcal{M}_{f}(\mathbb{N})$. By Poincaré Recurrence Theorem we have that $\mu$-almost every point $n\in\mathbb{N}$ is recurrent. Given $n_{0}\in\mathbb{N}$ a recurrent point, for every open set $n_{0}\in\mathcal{U}$ we have $f^{k}(n_{0})\in\mathcal{U}$ for some $k\in\mathbb{N}$. Hence, by taking $\mathcal{U}=\{n_{0},2n_{0}\}$ we have $f^{k}(n_{0})\in\{n_{0},2n_{0}\}$ for some $k\in\mathbb{N}$. It means that either $f^{k}(n_{0})=n_{0}$ or $f^{k}(n_{0})=2n_{0}$, that is, $f^{k+1}(n_{0})=n_{0}$. ∎
3.3. The coarseness of the key topology
The next lemma shows the necessity of working without continuity and still obtain periodicity.
Lemma 3.
The topology $\mathcal{T}$ given in Lemma 2 is coarser than the discrete one.
Proof.
[…]
3.4. Peculiarity of the discrete topology
The following lemma shows the peculiarity of the discrete topology by making the map continuous.
Lemma 4.
The unique topology making the Collatz function $f$ continuous and containing the collection $\{\{n,2n\}\mid n\in\mathbb{N}\}$ is the discrete one.
Proof.
It is enough to show that if a topology satisfies the hypothesis then it contains all the singletons. In fact, we first show that it contains the even singletons $\{2k\}$. Once it contains $\{k,2k\}$ and $\{2k,4k\}$ it must contain the intersection $\{2k\}=\{k,2k\}\cap\{2k,4k\}$. Now, we show that it contains the odd singletons. Given an odd singleton $\{n\}$ we have that $n\in f^{-1}(\{(3n+1)/2,3n+1\})=\{n,3n+1\}$ which is open because by hypothesis $f$ is continuous and $\{(3n+1)/2,3n+1\}$ is open, since $n$ is odd and $3n+1$ is even. So, the following intersection is open $\{n\}=\{n,2n\}\cap\{n,3n+1\}.$ Once the topology contains all the singletons, it coincides with the discrete one. ∎
3.5. Some key remarks
Remark 5.
For every periodic orbit there exists an ergodic $f$-invariant probability supported at the orbit and by Lemma 2 for every $f$-invariant probability there exists a periodic orbit. Therefore, the existence of periodic orbits is closely related to the existence of $f$-invariant probabilities.
Remark 6.
The topology $\mathcal{T}$ does not contain all the singletons since it does not coincide with the discrete one by Lemma 3. The orbit $\{1,2,4\}$ is open and then a Borelian. In fact, we have the following open sets:
$\{1,2,4\}=\{1,2\}\cup\{2,4\}\quad\text{is open}.$
Remark 7.
Given any ergodic probability $\mu\in\mathcal{M}_{f}(\mathbb{N})$ we have that $\mu$-almost every point in $\mathbb{N}$ is recurrent and by Lemma 2 they must be periodic. Then, $\mu$ is supported at a periodic orbit.
Remark 8.
Observe that we could take a recurrent point not related to any $f$-invariant probability a priori and conclude that it must be periodic. And then related to an ergodic probability.
Remark 9.
For every $x,y\in\mathbb{N}$ the following subset is open
$\{y,2y,4y,\cdots,2^{x}y\}.$
In fact, we have
$\{y,2y,4y,\cdots,2^{x}y\}=\{y,2y\}\cup\{2y,4y\}\cup\cdots\{2^{x-1}y,2^{x}y\}.$
As a consequence, every orbit of this type is an open subset. Also, there are orbits arbitrarily long. Also, every periodic orbit is an open subset.
3.6. Exploring periodicity towards integrals
Assume that there exists $\delta\in\mathcal{M}_{f}(\mathbb{N})$ ergodic such that $\delta\neq\delta_{0}$. Denote by $\mathcal{O}(x)$ the orbit of the point $x\in\mathbb{N}$. The probability $\delta$ is also supported on a periodic orbit $\mathcal{O}(x)$ for some $x\in\mathbb{N}$. We can compute for a potential $\phi$
$\int\phi d\delta=\frac{1}{\#\mathcal{O}(x)}\sum_{i=0}^{\#\mathcal{O}(x)-1}\phi(f^{i}(x)).$
3.7. A key potential
The following lemma is key in the building of the bridge. It construct a special continuous and unbounded potential that will be useful in the future.
Lemma 10.
There exists a $\mathcal{T}$-continuous potential $\varphi:\mathbb{N}\to\mathbb{N}$ which is constant on periodic orbits whose integral with respect to an ergodic probability $\delta$ supported on $\mathcal{O}(x)$ is given by
$\int\varphi d\delta=\sum_{i\in\mathcal{O}(x)}i.$
Proof.
Define $\varphi$ as
[…]
We obtain $\varphi:\mathbb{N}\to\mathbb{N}$ continuous with respect to $\mathcal{T}$ and
$\int\varphi d\delta=\frac{1}{\#\mathcal{O}(x)}\sum_{i=0}^{\#\mathcal{O}(x)-1}\varphi(f^{i}(x))=\frac{\#\mathcal{O}(x)\sum_{i\in\mathcal{O}(x)}i}{\#\mathcal{O}(x)}=\sum_{i\in\mathcal{O}(x)}i.$
In fact, we have $\varphi^{-1}(\{n,2n\}),n\neq 0$, pre-image of an open set in the base, either empty or union of periodic orbits. ∎
3.8. The ergodic decomposition
The next lemma guarantees that we can decompose any $f$-invariant measure into a convex sum of ergodic probabilities even in a noncompact space.
Lemma 11.
Any $f$-invariant probability $\mu$ is a convex combination of ergodic probabilities.
Proof.
This would be a consequence of the Ergodic Decomposition Theorem, but we cannot use it here because the topology is not even metrizable.
However, once the support of $\mu$ is a forward invariant subset, it must give full mass to the recurrent points, then giving full mass to the set of periodic orbits, being a union of periodic orbits almost surely. Then, we conclude that the measure $\mu$ is a convex combination of probabilities supported on periodic orbits, then ergodic ones. To be clear, let $\{\mathcal{O}_{i}\}$ be the countable collection of periodic orbits contained in the support of $\mu$. Once it is the subset of recurrent points and countable, we have
$\mu\Bigg(\bigcup_{i}\mathcal{O}_{i}\Bigg)=\sum_{i}\mu(\mathcal{O}_{i})=\sum_{i}\mu(\mathcal{O}_{i})\delta_{i}(\mathcal{O}_{i})=1.$
We conclude that $\mu$ is a convex combination of $\delta_{i}$ the ergodic probabilities supported at $\mathcal{O}_{i}$ because $\mu(\mathcal{O}_{i})\in[0,1],\delta_{i}(\mathcal{O}_{i})=1$ for every $i$. ∎
3.9. Finiteness of periodic orbits
The following lemma is an equivalence between finiteness of periodic orbits and the existence of equilibrium states.
Lemma 12.
Finiteness of periodic orbits is equivalent to every continuous potential $\phi:\mathbb{N}\to\mathbb{N}$ with respect to $\mathcal{T}$ possessing at least one equilibrium state.
Proof.
By definition, the pressure $P(\phi)$ for any measurable (in particular continuous) potential $\phi$ is given by the following supremum:
(1) $P(\phi)=\sup_{\mu\in\mathcal{M}_{f}(\mathbb{N})}\left\{h_{\mu}(f)+\int\phi\,d\mu\right\}=\sup_{\mu\in\mathcal{M}_{f}(\mathbb{N})}\left\{\int\phi\,d\mu\right\}$
where $h_{\mu}(f)=0$ for every $f$-invariant probability $\mu$ because each ergodic probability is supported on a periodic orbit, then having zero entropy. Also, any general $f$-invariant probability is convex combination of ergodic ones as proved in Lemma 11. By definition, an equilibrium state is a measure which attains the supremum.
If we have finitely many periodic orbits, then
$\sup_{\mu\in\mathcal{M}_{f}(\mathbb{N})}\left\{\int\phi\,d\mu\right\}<\infty.$
There exists $\delta$ such that
$\sup_{\mu\in\mathcal{M}_{f}(\mathbb{N})}\left\{\int\phi\,d\mu\right\}=\max_{\mu\in\mathcal{M}_{f}(\mathbb{N})}\left\{\int\phi\,d\mu\right\}=\int\phi d\delta.$
Conversely, the existence of an equilibrium state implies that for every continuous and unbounded potential $\phi$ we have
$\sup_{\mu\in\mathcal{M}_{f}(\mathbb{N})}\left\{\int\phi\,d\mu\right\}<\infty.$
Then, there must exist finitely many periodic orbits. Otherwise, by looking at the particular potential $\varphi$ defined above, the integral $\int\varphi d\delta=\sum_{i\in\mathcal{O}(x)}i$ can attains arbitrarily large values by continuity (and measurability) and unboundedness. ∎
Remark 13.
While the proof of the conjecture requires uniqueness of periodic orbits, we describe a mechanism from thermodynamic formalism to obtain finiteness of periodic orbits, which is equivalent to finiteness of ergodic probabilities.
3.10. Uniqueness of periodic orbits: the final step
The following lemma establishes a complete bridge between the Collatz Conjecture and the uniqueness of equilibrium states.
Lemma 14.
Uniqueness of periodic orbits is equivalent to every continuous and bounded potential $\phi:\mathbb{N}\to\mathbb{N}$ with respect to $\mathcal{T}$ possessing a unique equilibrium state.
Proof.
By definition, the pressure $P(\phi)$ for any measurable (in particular continuous) potential $\phi$ is given by the following supremum:
(2) $P(\phi)=\sup_{\mu\in\mathcal{M}_{f}(\mathbb{N})}\left\{h_{\mu}(f)+\int\phi\,d\mu\right\}=\sup_{\mu\in\mathcal{M}_{f}(\mathbb{N})}\left\{\int\phi\,d\mu\right\}$
where $h_{\mu}(f)=0$ for every $f$-invariant probability $\mu$ because each ergodic probability is supported on a periodic orbit, then having zero entropy. Also, any general $f$-invariant probability is convex combination of ergodic ones by Lemma 11. If we have a unique periodic orbit $\{1,2,4\}$ with the unique ergodic probability $\delta_{0}$, then readily
$\sup_{\mu\in\mathcal{M}_{f}(\mathbb{N})}\left\{\int\phi d\mu\right\}=\int\phi d\delta_{0}$
and $\delta_{0}$ is the unique equilibrium state for any continuous potential $\phi$. By definition, an equilibrium state is a measure which attains the supremum.
Conversely, the existence of a unique equilibrium state for any continuous potential $\phi$ implies the uniqueness of periodic orbits as follows. Set
$\mathcal{O}:=\bigcup_{x\,\,\text{is periodic}}\mathcal{O}(x)$
Denoting by $\chi_{X}$ the characteristic function of $X\subset\mathbb{N}$, we have that any ergodic probability $\delta$, which is supported on a periodic orbit is an equilibrium state because
$1=\int\chi_{\mathcal{O}}d\delta=\sup_{\mu\in\mathcal{M}_{f}(\mathbb{N})}\left\{\int\chi_{\mathcal{O}}d\mu\right\}=P(\chi_{\mathcal{O}}).$
Once by hypothesis there exists a unique equilibrium state for $\chi_{\mathcal{O}}$ (which is bounded and continuous because every orbit is an open subset), we conclude that there exists a unique ergodic probability $\delta_{0}$ and the unique periodic orbit is $\{1,2,4\}$. ∎
Remark 15.
While the characteristic function $\chi_{X}$ is measurable in general, in our $\sigma$-algebra it depends on the subset $X\subset\mathbb{N}$. In the particular case of $X=\mathcal{O}$, once the periodic orbits are open subsets, we have that $\chi_{\mathcal{O}}$ is continuous. Moreover, in Lemma 14 we address an alternative approach of the conjecture, rather than a proof of it.
4. Finiteness of Cycles
In this section, we use the Alexandroff Compactification of topology $\mathcal{T}$ to prove finiteness of cycles. We state it as a theorem.
Theorem 16.
Consider the Collatz map $f:\mathbb{N}\to\mathbb{N}$ and the topology $\mathcal{T}$ and the Borel $\sigma$-algebra $\Sigma$ given in Lemma 2. By using the Alexandroff Compactification of $\mathcal{T}$, we obtain finiteness of cycles for $f$.
Proof.
Recalling the process of compactification by Alexandroff, we know that the open sets in the compactification $\hat{\mathcal{T}}$ of the set $\hat{\mathbb{N}}$$\mathbb{N}\cup\{\infty\}$ is given by the open sets $U\in\mathcal{T}$ together with sets of the form $V=(\mathbb{N}\backslash K)\cup\{\infty\}$, where $K$ is a compact set in the topology $\mathcal{T}$. We observe that the extended map $\hat{f}$$\hat{\mathbb{N}}\to\hat{\mathbb{N}}$ is defined by
[…]
Assuming, by contradiction, that there are infinitely many cycle $\mathcal{O}_{1},\cdots,\mathcal{O}_{n},\cdots$, we consider an ergodic measure $\delta_{i}$ supported at $\mathcal{O}_{i}$ and take any convex combination
$\mu:=\sum_{i}^{\infty}a_{i}\delta_{i}\quad\text{where}\quad\sum_{i}^{\infty}a_{i}=1\quad\text{and}\quad a_{i}>0.$
As we can see, the fixed point $n=\infty$ for $\hat{f}$ admits an ergodic $\delta_{\infty}$ measure supported at it, but it is distinct from the original ergodic measures $\delta_{i}$.
We claim that the support of $\mu$ is exactly the union of all cycles $\{\mathcal{O}_{i}\}_{i\geq 1}$. In fact, if we had a point $x\in\text{supp}\mu$ which is not in this union, considering the open set $\{x,2x\}$ we have $\mu(\{x,2x\})>0$, which implies that $\delta_{j}(\{x,2x\})>0$ for some $j\geq 1$. It means that $\{x,2x\}\cap\mathcal{O}_{j}\neq\emptyset$, which implies that $2x\in\mathcal{O}_{j}$, that is, $f(2x)=x\in\mathcal{O}_{j}$, which is a contradiction. Then,
$\text{supp}\mu=\bigcup_{i}^{\infty}\mathcal{O}_{i}$
Once every cycle is open, it means that $\text{supp}\mu$ is open. Also, as a consequence of the definition of support, we have that $\text{supp}\mu$ is closed. So, $\text{supp}\mu$ is a clopen subset in the topology $\mathcal{T}$.
Since the open and closed subsets of $\mathcal{T}$ are also open and closed subsets in the compactification $\hat{\mathcal{T}}$, we conclude that $\text{supp}\mu$ is clopen in $\hat{\mathcal{T}}$. As every closed subset of a compact set is also compact, the subset $\text{supp}\mu$ is compact, having the union $\bigcup_{i}^{\infty}\mathcal{O}_{i}$ as a covering. By compactness, we can obtain a subcover $\bigcup_{j}^{p}\mathcal{O}_{i_{j}}$. Once the cycles are pairwise disjoint, we conclude that
$\text{supp}\mu=\bigcup_{i}^{\infty}\mathcal{O}_{i}=\bigcup_{j}^{p}\mathcal{O}_{i_{j}}.$
It means that there are at most finitely many cycles, as we wished. The theorem is proved. ∎
5. Uniqueness of Cycle
Theorem 17.
Consider the Collatz map $f:\mathbb{N}\to\mathbb{N}$ and the topology $\mathcal{T}$ and the Borel $\sigma$-algebra $\Sigma$ given in Lemma 2. The unique possible cycle is $\{1,2,4\}$.
Proof.
Let $\mathcal{O}_{1},\cdots,\mathcal{O}_{p}$ the finite list of cycles guaranteed by Theorem 16. Let us assume, by contradiction, then there exists $\mathcal{O}_{p}\neq\{1,2,4\}=\mathcal{O}_{1}$. As a consequence of the proof we know that the following set is clopen, compact and finite, with respect to the topology $\mathcal{T}$: $\mathcal{O}=\bigcup_{i=1}^{p}\mathcal{O}_{i}$.
Let $x:=\max\mathcal{O}$, which must be even because otherwise we would have $x<3x+1\in\mathcal{O}$. So $\{x,x/2\}\subset\mathcal{O}$. We have $x/2$ even because otherwise we would have $x<3(x/2)+1\in\mathcal{O}$. Also, there must exist an odd number $y\in\mathcal{O}$ such that $3y+1=x$ because otherwise $x<2x\in\mathcal{O}$. But we have the following geometric progression contained in $\mathcal{O}$: $\{y,2y,\cdots,2^{k}y\}$. But since $4y>3y+1=x=\max\mathcal{O}$, we must have $k=1$ and $\{y,2y\}\subset\mathcal{O}$. So, the following numbers are in the set $\mathcal{O}$:
$2y,y,x=3y+1,\frac{x}{2}=\frac{3y+1}{2}=y+\frac{y+1}{2}<y+y=2y.$
But we have
$y<\frac{3y+1}{2}<2y$
and the number $\frac{x}{2}=\frac{3y+1}{2}$ would be an even number in $\mathcal{O}$ between $y$ and $2y$. We also can find $z\in\mathcal{O}$ and odd number such that $3z+1=2y$. Also, the sequence $z,2z,\cdots,2^{m}z\in\mathcal{O}$. But since $8z>6z+2=2(3z+1)=2\cdot 2y=4y>x$ we must have $m=2$ and $z,2z,4z\in\mathcal{O}$. Moreover, we can find $w\in\mathcal{O}$ an odd number such that $3w+1=4z$ and $w,2w,\cdots,2^{q}w\in\mathcal{O}$. But $8w>6w+2=2(3w+1)=2\cdot 4z=8z>x$. So, we must have $q=2$ and $w,2w,4w\in\mathcal{O}$. We can find $t\in\mathcal{O}$ an odd number such that $3t+1=4w$. So $8t>6t+2=2(3t+1)=2\cdot 4w=8w>x$. By induction, we must find $a\in\mathcal{O}$ odd such that $a,2a,4a=x$. We have that
$z<y,z<w<t<\cdots<a.$
Since the cycles are pairwise disjoint, we can suppose that $\mathcal{O}_{p}$ is the unique cycle such that $x,y,z,w,t,a\in\mathcal{O}_{p}$. Moreover, the odd number $z=\min\mathcal{O}_{p}$. Then the orbit is
$\cdots\to\min\mathcal{O}_{p}=z\to 3z+1=2y\to y\to 3y+1=x=\max\mathcal{O}_{p}\to\cdots$
Then,
$x=3\frac{3z+1}{2}+1\implies x=\frac{9z+3}{2}+1\implies\max\mathcal{O}_{p}=\frac{9\min\mathcal{O}_{p}+3}{2}+1$
Also, we have $r$ geometric progressions with $3$ elements as $x=4a,2a,a$ and $1$ pair $2y,y$ and then the cycle $\mathcal{O}_{p}$ has $3r+2$ elements. We have an orbit as follows
$x=4a,2a,a,3a+1=4b,2b,b,3b+1=4c,2c,c.\cdots,4z,2z,z,3z+1=2y,y,3y+1=x.$
So, 4a4b=4a(3a+1)=a14a-4b=4a-(3a+1)=a-1, 4b4c=b1,4w4z=w14b-4c=b-1,4w-4z=w-1, which implies
$x=4a-4b+4b-4c+4c-\cdots+4w-4z+4z=$
$a-1+b-1+c-1+\cdots+w-1+z+3z=a+b+c+\cdots w+z-r+2y-1=$
$a+b+c+\cdots w+z+y-r+y-1=\frac{9z+3}{2}+1.$
The list $a,b,c,\cdots,w,z,y$ has $r+1$ odd numbers. If $r$ is odd, then $r+1$ is even and the sum $a+b+c+\cdots+w+z+y$ is even. So, yr1y-r-1 is odd and we get a contradiction because $x$ is even and the following sum is odd
$a+b+c+\cdots w+z+y-r+y-1.$
Then, $r$ must be even and we have that the sum $a+b+c+\cdots+w+z+y$ is odd. But yr1y-r-1 is even and the following sum is odd
$a+b+c+\cdots w+z+y-r+y-1.$
This contradiction shows that the cycle $\mathcal{O}_{p}$ cannot exist. We conclude that the unique cycle is $\mathcal{O}=\mathcal{O}_{1}=\{1,2,4\}$. ∎
6. Baker and Syracuse Maps
In this section we show that the same technique works for the Baker map and, more generally, for the so-called Syracuse maps, that is, for $a,b\in\mathbb{Z}$ where $a,b$ are odd numbers, we define the Syracuse map as
[…]
The Collatz map is a Syracuse one for $a=3,b=1$ and the Baker map is the one for a=3,b=1a=3,b=-1. We observe that in the construction of our topology and $\sigma$-algebra the part of the map assigning values for odd numbers plays no fundamental role in the results obtained. It means that the same technique can be applied to obtain analogous key topology and $\sigma$-algebra for the general case. We then obtain, analogously, that the number of cycles is finite for any Syracuse map. For example, there are only three known cycles for the Baker map:
$1\to 2\to 1,\quad 5\to 14\to 7\to 20\to 10\to 5$
and
$17\to 50\to 25\to 74\to 37\to 110\to 55\to 164\to 82\to 41\to$
$122\to 61\to 182\to 91\to 272\to 136\to 68\to 34\to 17.$
Denoting by $\mathcal{O}_{1},\mathcal{O}_{2},\mathcal{O}_{3}$ the cycles and by $\mathcal{O}=\mathcal{O}_{1}\cup\mathcal{O}_{2}\cup\mathcal{O}_{3}$, we have
$1=\int\chi_{\mathcal{O}}d\delta=\sup_{\mu\in\mathcal{M}_{f}(\mathbb{N})}\left\{\int\chi_{\mathcal{O}}d\mu\right\}=P(\chi_{\mathcal{O}}).$
Also, taking $\delta_{i}$ the ergodic measure supported at $\mathcal{O}_{i}$ we have
$1=\int\chi_{\mathcal{O}}d\delta_{i}=\sup_{\mu\in\mathcal{M}_{f}(\mathbb{N})}\left\{\int\chi_{\mathcal{O}}d\mu\right\}=P(\chi_{\mathcal{O}}).$
So, each measure $\delta_{i}$ is an equilibrium state. The lack of uniqueness of equilibrium states is coherent with the lack of uniqueness of cycles. It is reasonable to expect that the Collatz map has uniqueness of equilibrium state. The role of the part $an+b$ seems to be related to the quantity of cycles for the map.
7. Divergent Orbits
Theorem 18.
Consider the Collatz map $f:\mathbb{N}\to\mathbb{N}$ and the topology $\mathcal{T}$ and the Borel $\sigma$-algebra $\Sigma$ given in Lemma 2. There is no divergent orbits.
Proof.
In fact, for every full orbit $\{f^{k}(n)\}_{k\in\mathbb{Z}}$, it is a clopen subset because it is a union of geometric progressions with ration $2$ and the forward orbit $\{f^{k}(n)\}_{k\in\mathbb{N}}$ is a union of finite geometric progressions with ration $2$. Its complement is also like this. We can take it as a covering. Since it is closed, it gets compact after the Alexandroff Compactification made in Section 4. Taking the union of geometric progressions itself as a covering, we can take a finite subcovering. So, the forward orbit $\{f^{k}(n)\}_{k\in\mathbb{N}}$ is a finite union of finite geometric progressions with ration $2$. It shows that the forward orbit is bounded and it must enter a cycle. Since the unique cycle is $\{1,2,4\}$ we must have every orbit reaching $1$. ∎
Remark 19.
We observe that the full orbit $\{f^{k}(n)\}_{k\in\mathbb{Z}}$ may contain and infinite geometric progression with ration $2$ but the forward orbit $\{f^{k}(n)\}_{k\in\mathbb{N}}$ cannot.