### v4

# On The Collatz Conjecture: Topological and Ergodic Approach

Abstract.
We study a class of maps having the Collatz function (famously related to the Collatz Conjecture) as an example, under the topological and ergodic perspectives, including an approach with thermodynamic formalism. By introducing a key topology and its Borel $\sigma$-algebra we show that recurrence implies periodicity. Moreover, we establish that the set of periodic orbits is finite if, and only if, every continuous potential possesses some equilibrium state. The uniqueness of periodic orbits is equivalent to the uniqueness of equilibrium state for every bounded and continuous potential. Additionally, by using the dictionary established in the paper, we prove finiteness of cycles, which is a significant advance to the conjecture itself. Finally, we apply our technique to the Baker and Syracuse maps, obtaining a similar result on the finiteness of orbits for a general class of important maps.
1. Introduction
We study a family of maps which generalizes the Collatz map, defined into two cases, namely when the natural number is either even or odd. For the case where the number $n$ is even, we always take it into its half $\frac{n}{2}$. We develop the topological and ergodic approach of this family of maps to obtain a result regarding the quantity of periodic orbits, establishing that for any of them it is finite. The quantity would be probably established by using the odd case.
Given any function $f_{0}:\mathbb{N}\to\mathbb{N}$ we can define another function $f:\mathbb{N}\to\mathbb{N}$ as follows:
[…]
As an example, we have the famous Collatz function $f:\mathbb{N}\to\mathbb{N}$ defined for $f_{0}(n)=3n+1$ as follows:
[…]
It has been extensively studied because of the famous Collatz Conjecture: For each $n\in\mathbb{N}$ there exists $k\in\mathbb{N}$ such that $f^{k}(n)=1$. In other words, every orbit enters the cycle $\{1,2,4\}$. This is the unique cycle known.
We intend to obtain some results on the topological and ergodic aspects of the Collatz function. It will be useful in the study of the Collatz Conjecture. We highlight its difficult by building a bridge between it and the existence and uniqueness of equilibrium states for every continuous potentials with respect to a key topology. It provides a ”dictionary” which perhaps turns possible the complete proof in another realm. We provide a significant advance by proving finiteness of cycles. To prove the conjecture, it remains to prove that there exists a unique cycle and no divergent orbits. We also apply our technique for the Baker and Syracuse maps as follows where $a,b\in\mathbb{N}$ are odd numbers:
[…]
2. Main result
The following result is an attempt to link a very hard and famous problem on Number Theory to Ergodic Theory in the realm of Themodynamic Formalism. We list some topological and ergodic properties of the family of functions defined above which highlight both the beauty and the difficulty of the Collatz Conjecture using the language of equilibrium states.
Theorem A.
We have the following topological and ergodic facts with respect to $f$
  • There exists a topology coarser than the discrete one and a $\sigma$-algebra with respect to which every recurrent point is periodic and the ergodic probabilities are the ones supported on the periodic orbits.
  • If $f$ the Collatz map, then it is measurable with respect to this $\sigma$-algebra but not continuous, once $f$ is only continuous with respect to the discrete one.
  • Every periodic orbit is an open subset.
  • Every $f$-invariant probability is a convex linear combination of ergodic probabilities and has zero entropy.
  • Finiteness of periodic orbits is guaranteed by the existence of an equilibrium state for every continuous potential $\phi:\mathbb{N}\to\mathbb{N}$ which is integrable with respect to any $f$-invariant probability.
  • Uniqueness of periodic orbits is guaranteed by the uniqueness of equilibrium state for every bounded and continuous potential $\phi:\mathbb{N}\to\mathbb{N}$.
By using Theorem A we can prove the following theorem for the general class of maps $f$, incluing the Collatz, Bake and Syracuse maps.
Theorem B.
For the family of maps defined in ()(*) there exists at most finitely many periodic orbits. Moreover, it may exists no periodic orbit.
3. Proof of Theorem A
We divide the proof of our Main Theorem into some lemmas and remarks. We proceed with this now.
3.1. Invoking foundations
The first lemma invokes foundations of General Topology and Measure Theory to build both a key topology and a key $\sigma$-algebra.
Lemma 1.
Let $\mathcal{T}_{\lambda}$ be a family of topologies on $\mathbb{N}$ and $\sigma(\mathcal{T}_{\lambda})$ their Borel $\sigma$-algebras. We have that
$\mathcal{T}:=\bigcap_{\lambda}\mathcal{T}_{\lambda}\quad\text{and}\quad\Sigma:=\bigcap_{\lambda}\sigma(\mathcal{T}_{\lambda})$
are respectively a topology and a $\sigma$-algebra on $\mathbb{N}$ if a map $f:\mathbb{N}\to\mathbb{N}$ is measurable with respect to every $\sigma(\mathcal{T}_{\lambda})$ it does with respect to $\Sigma$.
Proof.
It is well known from the theory of General Topology and Measure Theory that $\mathcal{T}$ and $\Sigma$ are respectively a topology and a $\sigma$-algebra. Moreover, given any $A\in\Sigma$ we have $A\in\sigma(T_{\lambda})$ for every $\lambda$, which implies $f^{-1}(A)\in\sigma(T_{\lambda})$ for every $\lambda$ because $f$ is measurable with respect to $\sigma(\mathcal{T}_{\lambda})$ and we conclude that $f^{-1}(A)\in\Sigma$. Since $A$ is arbitrary, we obtain $f$ measurable with respect to $\Sigma$. ∎
3.2. Trapping the orbits towards periodicity
The following lemma builds both a topology and its Borel $\sigma$-algebra making the Collatz map both measurable and predictable in the language of Ergodic Theory.
Lemma 2.
There exist a topology $\mathcal{T}$ and a Borel $\sigma$-algebra $\Sigma$ with respect to which the Collatz map $f:\mathbb{N}\to\mathbb{N}$ possesses $f$-invariant Borel probabilities. Moreover, if there exists some $f$-invariant probability, then every recurrent point is periodic.
Proof.
Endow $\mathbb{N}$ with $\mathcal{T}$ as the intersection of all topologies (the coarsest) containing the following collection of subsets
$\{\{n,2n\}\mid n\in\mathbb{N}\}.$
By Lemma 1 $\mathcal{T}$ is well defined and we consider the $\sigma$-algebra $\Sigma$ constructed also as in Lemma 1. We obtain $f$ measurable with respect to $\Sigma$.
Let $\mathcal{M}_{f}(\mathbb{N})$ be the set of all Borelian $f$-invariant probabilities. Once there exists at least one $f$-invariant probability, we can take $\mu\in\mathcal{M}_{f}(\mathbb{N})$.
By Poincaré Recurrence Theorem we have that $\mu$-almost every point $n\in\mathbb{N}$ is recurrent. Given $n_{0}\in\mathbb{N}$ a recurrent point, for every open set $n_{0}\in\mathcal{U}$ we have $f^{k}(n_{0})\in\mathcal{U}$ for some $k\in\mathbb{N}$. Hence, by taking $\mathcal{U}=\{n_{0},2n_{0}\}$ we have $f^{k}(n_{0})\in\{n_{0},2n_{0}\}$ for some $k\in\mathbb{N}$. It means that either $f^{k}(n_{0})=n_{0}$ or $f^{k}(n_{0})=2n_{0}$, that is, $f^{k+1}(n_{0})=n_{0}$. ∎
Remark 3.
Once there exists some periodic orbit, we can consider some ergodic probability $\delta_{0}\in\mathcal{M}_{f}(\mathbb{N})$.
3.3. The coarseness of the key topology
The next lemma shows the necessity of working without continuity and still obtain periodicity.
Lemma 4.
The topology $\mathcal{T}$ given in Lemma 2 is coarser than the discrete one.
Proof.
[…]
3.4. Peculiarity of the discrete topology for Collatz
The following lemma shows the peculiarity of the discrete topology by making the map continuous.
Lemma 5.
The unique topology making the Collatz function $f$ continuous and containing the collection $\{\{n,2n\}\mid n\in\mathbb{N}\}$ is the discrete one.
Proof.
It is enough to show that if a topology satisfies the hypothesis then it contains all the singletons. In fact, we first show that it contains the even singletons $\{2k\}$. Once it contains $\{k,2k\}$ and $\{2k,4k\}$ it must contain the intersection $\{2k\}=\{k,2k\}\cap\{2k,4k\}$. Now, we show that it contains the odd singletons. Given an odd singleton $\{n\}$ we have that $n\in f^{-1}(\{(3n+1)/2,3n+1\})=\{n,3n+1\}$ which is open because by hypothesis $f$ is continuous and $\{(3n+1)/2,3n+1\}$ is open, since $n$ is odd and $3n+1$ is even. So, the following intersection is open $\{n\}=\{n,2n\}\cap\{n,3n+1\}.$ Once the topology contains all the singletons, it coincides with the discrete one. ∎
3.5. Some key remarks
Remark 6.
For every periodic orbit there exists an ergodic $f$-invariant probability supported at the orbit and by Lemma 2 for every $f$-invariant probability there exists a periodic orbit. Therefore, the existence of periodic orbits is closely related to the existence of $f$-invariant probabilities.
Remark 7.
Given any ergodic probability $\mu\in\mathcal{M}_{f}(\mathbb{N})$ we have that $\mu$-almost every point in $\mathbb{N}$ is recurrent and by Lemma 2 they must be periodic. Then, $\mu$ is supported at a periodic orbit.
Remark 8.
Observe that we could take a recurrent point not related to any $f$-invariant probability a priori and conclude that it must be periodic. And then related to an ergodic probability.
Remark 9.
For every $x,y\in\mathbb{N}$ the following subset is open
$\{y,2y,4y,\cdots,2^{x}y\}.$
In fact, we have
$\{y,2y,4y,\cdots,2^{x}y\}=\{y,2y\}\cup\{2y,4y\}\cup\cdots\{2^{x-1}y,2^{x}y\}.$
As a consequence, every orbit of this type is an open subset. Also, there are orbits arbitrarily long. Also, every periodic orbit is an open subset.
3.6. Exploring periodicity towards integrals
Assume that there exists $\delta\in\mathcal{M}_{f}(\mathbb{N})$ ergodic such that $\delta\neq\delta_{0}$. Denote by $\mathcal{O}(x)$ the orbit of the point $x\in\mathbb{N}$. The probability $\delta$ is also supported on a periodic orbit $\mathcal{O}(x)$ for some $x\in\mathbb{N}$. We can compute for a potential $\phi$
$\int\phi d\delta=\frac{1}{\#\mathcal{O}(x)}\sum_{i=0}^{\#\mathcal{O}(x)-1}\phi(f^{i}(x)).$
3.7. A key potential
The following lemma is key in the building of the bridge. It construct a special continuous and unbounded potential that will be useful in the future.
Lemma 10.
There exists a $\mathcal{T}$-continuous potential $\varphi:\mathbb{N}\to\mathbb{N}$ which is constant on periodic orbits whose integral with respect to an ergodic probability $\delta$ supported on $\mathcal{O}(x)$ is given by
$\int\varphi d\delta=\sum_{i\in\mathcal{O}(x)}i.$
Proof.
Define $\varphi$ as
[…]
We obtain $\varphi:\mathbb{N}\to\mathbb{N}$ continuous with respect to $\mathcal{T}$ and
$\int\varphi d\delta=\frac{1}{\#\mathcal{O}(x)}\sum_{i=0}^{\#\mathcal{O}(x)-1}\varphi(f^{i}(x))=\frac{\#\mathcal{O}(x)\sum_{i\in\mathcal{O}(x)}i}{\#\mathcal{O}(x)}=\sum_{i\in\mathcal{O}(x)}i.$
In fact, we have $\varphi^{-1}(\{n,2n\}),n\neq 0$, pre-image of an open set in the base, either empty or union of periodic orbits. ∎
3.8. The ergodic decomposition
The next lemma guarantees that we can decompose any $f$-invariant measure into a convex sum of ergodic probabilities even in a noncompact space.
Lemma 11.
Any $f$-invariant probability $\mu$ is a convex combination of ergodic probabilities.
Proof.
This would be a consequence of the Ergodic Decomposition Theorem, but we cannot use it here because the topology is not even metrizable.
However, once the support of $\mu$ is a forward invariant subset, it must give full mass to the recurrent points, then giving full mass to the set of periodic orbits, being a union of periodic orbits almost surely. Then, we conclude that the measure $\mu$ is a convex combination of probabilities supported on periodic orbits, then ergodic ones. To be clear, let $\{\mathcal{O}_{i}\}$ be the countable collection of periodic orbits contained in the support of $\mu$. Once it is the subset of recurrent points and countable, we have
$\mu\Bigg(\bigcup_{i=1}^{\infty}\mathcal{O}_{i}\Bigg)=\sum_{i=1}^{\infty}\mu(\mathcal{O}_{i})=\sum_{i=1}^{\infty}\mu(\mathcal{O}_{i})\delta_{i}(\mathcal{O}_{i})=1.$
We conclude that $\mu$ is a convex combination of $\delta_{i}$ the ergodic probabilities supported at $\mathcal{O}_{i}$ because $\mu(\mathcal{O}_{i})\in[0,1],\delta_{i}(\mathcal{O}_{i})=1$ for every $i$. ∎
3.9. The entropy as a convex combination
We prove that the entropy of a measure is a convex combination of the entropies of the ergodic measures, as a consequence of the ergodic decomposition.
Lemma 12.
Given
$\mu=\sum_{i=1}^{\infty}a_{i}\delta_{i},\sum_{i=1}^{\infty}a_{i}=1,$
we have
$h_{\mu}(f)=\sum_{i=1}^{\infty}a_{i}h_{\delta_{i}}(f).$
Proof.
It holds that
$\mu=\sum_{i=1}^{\infty}a_{i}\delta_{i}=\sum_{i=1}^{k}a_{i}\delta_{i}+\sum_{i=k+1}^{\infty}a_{i}\delta_{i}=\sum_{i=1}^{k}a_{i}\cdot\sum_{i=1}^{k}\frac{a_{i}}{\sum_{i=1}^{k}a_{i}}\delta_{i}+\sum_{i=k+1}^{\infty}a_{i}\cdot\sum_{i=k+1}^{\infty}\frac{a_{i}}{\sum_{i=k+1}^{\infty}a_{i}}\delta_{i}.$
Writing
$t_{k}=\sum_{i=k+1}^{\infty}a_{i},\mu_{k}=\sum_{i=1}^{k}\frac{a_{i}}{\sum_{i=1}^{k}a_{i}}\delta_{i},\nu_{k}=\sum_{i=k+1}^{\infty}\frac{a_{i}}{\sum_{i=k+1}^{\infty}a_{i}}\delta_{i},$
we obtain
$h_{\mu}(f)=h_{(1-t_{k})\mu_{k}+t_{k}\nu_{k}}(f)=(1-t_{k})h_{\mu_{k}}(f)+t_{k}h_{\nu_{k}}(f)=\sum_{i=1}^{k}a_{i}\cdot h_{\mu_{k}}(f)+t_{k}h_{\nu_{k}}(f)=$
$\sum_{i=1}^{k}a_{i}\cdot\sum_{i=1}^{k}\frac{a_{i}}{\sum_{i=1}^{k}a_{i}}h_{\delta_{i}(f)}+t_{k}h_{\nu_{k}}(f)=\sum_{i=1}^{k}a_{i}h_{\delta_{i}}(f)+t_{k}h_{\nu_{k}}(f).$
Taking $k\to\infty$, we obtain $t_{k}\to 0$ and
$h_{\mu}(f)=\sum_{i=1}^{\infty}a_{i}h_{\delta_{i}}(f).$
Remark 13.
Once we have $h_{\delta_{i}}(f)=0$ for every $i\geq 1$, then we obtain $h_{\mu}(f)=0$ for every $f$-invariant measure $\mu$.
3.10. Finiteness of periodic orbits
The following lemma is an equivalence between finiteness of periodic orbits and the existence of equilibrium states.
Lemma 14.
Finiteness of periodic orbits is equivalent to every continuous potential $\phi:\mathbb{N}\to\mathbb{N}$ with respect to $\mathcal{T}$ and integrable with respect to any $f$-invariant probability possessing at least one equilibrium state.
Proof.
By definition, the pressure $P(\phi)$ for any measurable (in particular continuous) potential $\phi$ is given by the following supremum:
(1) $P(\phi)=\sup_{\mu\in\mathcal{M}_{f}(\mathbb{N})}\left\{h_{\mu}(f)+\int\phi\,d\mu\right\}=\sup_{\mu\in\mathcal{M}_{f}(\mathbb{N})}\left\{\int\phi\,d\mu\right\}$
where $h_{\mu}(f)=0$ for every $f$-invariant probability $\mu$ because each ergodic probability is supported on a periodic orbit, then having zero entropy. Also, any general $f$-invariant probability is convex combination of ergodic ones as proved in Lemma 11. By definition, an equilibrium state is a measure which attains the supremum.
If we have finitely many periodic orbits, then
$\sup_{\mu\in\mathcal{M}_{f}(\mathbb{N})}\left\{\int\phi\,d\mu\right\}<\infty.$
There exists $\delta$ such that
$\sup_{\mu\in\mathcal{M}_{f}(\mathbb{N})}\left\{\int\phi\,d\mu\right\}=\max_{\mu\in\mathcal{M}_{f}(\mathbb{N})}\left\{\int\phi\,d\mu\right\}=\int\phi d\delta.$
Conversely, the existence of an equilibrium state implies that for every continuous and unbounded potential $\phi$ we have
$\sup_{\mu\in\mathcal{M}_{f}(\mathbb{N})}\left\{\int\phi\,d\mu\right\}<\infty.$
Then, there must exist finitely many periodic orbits. Otherwise, we could have $\int\phi d\delta_{i}\to\infty$ for $\phi$ unbounded on the support. ∎
Remark 15.
While the proof of the conjecture requires uniqueness of periodic orbits, we describe a mechanism from thermodynamic formalism to obtain finiteness of periodic orbits, which is equivalent to finiteness of ergodic probabilities.
3.11. Uniqueness of periodic orbits: the final step
The following lemma establishes a complete bridge between the Collatz Conjecture and the uniqueness of equilibrium states.
Lemma 16.
Uniqueness of periodic orbits is equivalent to every continuous and bounded potential $\phi:\mathbb{N}\to\mathbb{N}$ with respect to $\mathcal{T}$ possessing a unique equilibrium state.
Proof.
By definition, the pressure $P(\phi)$ for any measurable (in particular continuous) potential $\phi$ is given by the following supremum:
(2) $P(\phi)=\sup_{\mu\in\mathcal{M}_{f}(\mathbb{N})}\left\{h_{\mu}(f)+\int\phi\,d\mu\right\}=\sup_{\mu\in\mathcal{M}_{f}(\mathbb{N})}\left\{\int\phi\,d\mu\right\}$
where $h_{\mu}(f)=0$ for every $f$-invariant probability $\mu$ because each ergodic probability is supported on a periodic orbit, then having zero entropy. Also, any general $f$-invariant probability is convex combination of ergodic ones by Lemma 11. If we have a unique periodic orbit $\{1,2,4\}$ with the unique ergodic probability $\delta_{0}$, then readily
$\sup_{\mu\in\mathcal{M}_{f}(\mathbb{N})}\left\{\int\phi d\mu\right\}=\int\phi d\delta_{0}$
and $\delta_{0}$ is the unique equilibrium state for any continuous potential $\phi$. By definition, an equilibrium state is a measure which attains the supremum.
Conversely, the existence of a unique equilibrium state for any continuous potential $\phi$ implies the uniqueness of periodic orbits as follows. Set
$\mathcal{O}:=\bigcup_{x\,\,\text{is periodic}}\mathcal{O}(x)$
Denoting by $\chi_{X}$ the characteristic function of $X\subset\mathbb{N}$, we have that any ergodic probability $\delta$, which is supported on a periodic orbit is an equilibrium state because
$1=\int\chi_{\mathcal{O}}d\delta=\sup_{\mu\in\mathcal{M}_{f}(\mathbb{N})}\left\{\int\chi_{\mathcal{O}}d\mu\right\}=P(\chi_{\mathcal{O}}).$
Once by hypothesis there exists a unique equilibrium state for $\chi_{\mathcal{O}}$ (which is bounded and continuous because every orbit is an open subset), we conclude that there exists a unique ergodic probability $\delta_{0}$ and a unique periodic orbit. ∎
Remark 17.
While the characteristic function $\chi_{X}$ is measurable in general, in our $\sigma$-algebra it depends on the subset $X\subset\mathbb{N}$. In the particular case of $X=\mathcal{O}$, once the periodic orbits are open subsets, we have that $\chi_{\mathcal{O}}$ is continuous. Moreover, in Lemma 16 we address an alternative approach of the conjecture, rather than a proof of it.
4. Proof of Theorem B
In this section, we use the dictionary established in Theorem A to prove the existence of finitely many cycles. We state it as a theorem.
Theorem 18.
Consider the Collatz map $f:\mathbb{N}\to\mathbb{N}$ and the topology $\mathcal{T}$ and the Borel $\sigma$-algebra $\Sigma$ given in Lemma 2. By Lemma 14, we obtain the existence of at most finitely many cycles for $f$.
Proof.
Assuming, by contradiction, that there are infinitely many cycles $\mathcal{O}_{1},\cdots,\mathcal{O}_{k},\cdots$, for each $i\geq 1$ we consider the ergodic measure $\delta_{i}$ supported at $\mathcal{O}_{i}$ and take any convex combination
$\mu:=\sum_{i=1}^{\infty}a_{i}\delta_{i}\quad\text{where}\quad\sum_{i=1}^{\infty}a_{i}=1\quad\text{and}\quad a_{i}>0.$
We claim that the support of $\mu$ is exactly the union of all cycles $\{\mathcal{O}_{i}\}_{i\geq 1}$. In fact, if we had a point $x\in\text{supp}\mu$ which is not in this union, considering the open set $\{x,2x\}$ we have $\mu(\{x,2x\})>0$, which implies that $\delta_{j}(\{x,2x\})>0$ for some $j\geq 1$. It means that $\{x,2x\}\cap\mathcal{O}_{j}\neq\emptyset$, which implies that $2x\in\mathcal{O}_{j}$, that is, $f(2x)=x\in\mathcal{O}_{j}$, which is a contradiction. Then, $\text{supp}\mu=\bigcup_{i=1}^{\infty}\mathcal{O}_{i}$.
Given a continuous potential $\phi:\mathbb{N}\to\mathbb{N}$, integrable with respect to any $f$-invariant probability, we are going to show that there exists at least one equilibrium state.
$\sup_{\mu\in\mathcal{M}_{f}(\mathbb{N})}\left\{\int\phi d\mu\right\}=\int\phi d\delta_{i}$
In fact, we have
$\int\phi d\mu=\sum_{i=1}^{\infty}a_{i}\int\phi d\delta_{i}<\infty\implies a_{i}\int\phi d\delta_{i}\to 0.$
It is true for any convex combination and it implies that $\int\phi d\delta_{i}\leq M$ for some $M>0$. Otherwise, we could find a convex combination (integral) which diverges. In fact, if $\int\phi d\delta_{i}\to\infty$, we can take $\int\phi d\delta_{i_{j}}\geq 2^{j}$. Also, we can take $a_{i_{j}}=1/2^{j}$ and choose all the other $a_{k}$’s such that $\sum_{i=1}^{\infty}a_{i}=1$. Then,
$\int\phi d\mu=\sum_{i=1}^{\infty}a_{i}\int\phi d\delta_{i}\geq\sum_{j=1}a_{i_{j}}\int\phi d\delta_{i_{j}}\geq\sum_{j=1}^{\infty}(1/2^{j})\cdot 2^{j}=\infty.$
Once the values of the integrals are natural numbers, there must exist some $\delta_{i}$ such that it is an equilibrium state for $\phi$ by attaining the maximum integral. By Lemma 14, it means that there are finitely many cycles, as we wished. The theorem is proved. ∎
5. Baker and Syracuse Maps
In this section we observe that the same technique works for the Baker map and, more generally, for the so-called Syracuse maps, that is, for $a,b\in\mathbb{Z}$ where $a,b$ are odd numbers, we define the Syracuse map as
[…]
The Collatz map is a Syracuse one for $a=3,b=1$ and the Baker map is the one for a=3,b=1a=3,b=-1. We observe that in the construction of our topology and $\sigma$-algebra the part of the map assigning values for odd numbers plays no fundamental role in the results obtained. It means that the same technique can be applied to obtain analogous key topology and $\sigma$-algebra for the general case. We then obtain, analogously, that the number of cycles is finite for any Syracuse map. For example, there are only three known cycles for the Baker map:
$1\to 2\to 1,\quad 5\to 14\to 7\to 20\to 10\to 5$
and
$17\to 50\to 25\to 74\to 37\to 110\to 55\to 164\to 82\to 41\to$
$122\to 61\to 182\to 91\to 272\to 136\to 68\to 34\to 17.$
Denoting by $\mathcal{O}_{1},\mathcal{O}_{2},\mathcal{O}_{3}$ the cycles and by $\mathcal{O}=\mathcal{O}_{1}\cup\mathcal{O}_{2}\cup\mathcal{O}_{3}$, we have
$1=\int\chi_{\mathcal{O}}d\delta=\sup_{\mu\in\mathcal{M}_{f}(\mathbb{N})}\left\{\int\chi_{\mathcal{O}}d\mu\right\}=P(\chi_{\mathcal{O}}).$
Also, taking $\delta_{i}$ the ergodic measure supported at $\mathcal{O}_{i}$ we have
$1=\int\chi_{\mathcal{O}}d\delta_{i}=\sup_{\mu\in\mathcal{M}_{f}(\mathbb{N})}\left\{\int\chi_{\mathcal{O}}d\mu\right\}=P(\chi_{\mathcal{O}}).$
So, each measure $\delta_{i}$ is an equilibrium state. The lack of uniqueness of equilibrium states is coherent with the lack of uniqueness of cycles. It is reasonable to expect that the Collatz map has uniqueness of equilibrium state. The role of the part $an+b$ seems to be related to the quantity of cycles for the map.